OpenStudy (anonymous):

Use the Quadratic Formula to solve the equation. –x2 – x – 5 = 0

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

-x^2*

hero (hero):

\[–x^2 – x – 5 = 0\] 1st I would factor out -1: \[-(x^2 + x + 5) = 0\] Then divide both sides by -1 \[x^2 + x + 5 = 0\] Now I would use quadratic formula: \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] a = 1, b = 1, c = 5

OpenStudy (anonymous):

Oh, thank you very much :)

hero (hero):

Well, actually, hang on

hero (hero):

When in doubt, compute the discriminant to see if this is even has any real solution

hero (hero):

b^2 - 4ac = (1)^2 - 4(1)(5) = 1 - 20 = -19 No real solution

OpenStudy (anonymous):

Thats what i got also, i wanted to make sure i was doing it right. :)